# Bi-Level Online Provisioning and Scheduling with Switching Costs and Cross-Level Constraints

### v1

@@ V-C2 Value approximation error
The remaining terms capture the value difference between the true and surrogate objectives. By definition, $g_{k}(b)-\hat{g}_{k}(b)=\beta(L_{k}^{*}(b)-\hat{L}_{k}(b))$. Using the optimism lemma from the lower-level analysis (see [10]), we can show that $\hat{L}_{k}(b)$ (the value of the optimistic surrogate) is close to $L_{k}^{*}(b)$ (the true optimal value). The cumulative sum of these differences/errors is bounded by the lower-level regret bound, which thus yields the $\tilde{\mathcal{O}}(\sqrt{K})$ term in the regret. ∎
@@ V-D Final Regret Bound and Safety Guarantee
Theorem 1 establishes a near-optimal $\tilde{\mathcal{O}}(\sqrt{K})$ regret, implying the average regret vanishes. Importantly, the bound separates the roles of the two levels: the upper level contributes only logarithmic regret (and bounded switching overhead) thanks to strong convexity and stable updates, while the dominant $\tilde{\mathcal{O}}(\sqrt{K})$ term comes from learning the unknown lower-level dynamics safely even under time-varying budgets. This clean separation clarifies the benefit of our primal-dual interface. The extended LP provides a dual variable signal that converts lower-level learning progress into correct sensitivity feedback for upper-level subgradient in budget updating.
@@
Theorem 2 establishes the safety of our BALDE subroutine under dynamic/time-varying budget and constraint threshold. It guarantees that the lower-level scheduler strictly adheres to the resource limits imposed by the upper level.
@@
For the warm-up phase ($k\leq K_{0}$), safety is guaranteed by baseline policy $\pi_{\mathrm{base}}$. For $k>K_{0}$, applying the value difference lemma and Hölder’s inequality, the difference between the cost under true transition kernel $P$ and that under $P_{k}$ can be bound as $\left|V_{d}({\pi_{k}};P)-V_{d}({\pi_{k}};P_{k})\right|\leq\mathbb{E}[\sum_{t=1}^{T}\|P_{t}-P_{k,t}\|_{1}\|V_{d,t+1}^{\pi_{k}}\|_{\infty}]\leq TE[\sum_{t=1}^{T}\bar{\beta}_{t,k}^{p}|\pi_{k},P_{k}]=\epsilon_{k}^{\pi_{k}}(P_{k})$, where $V_{d,t}^{\pi}\triangleq\mathbb{E}[\sum\nolimits_{\tau=t}^{T}d_{t}(s_{t},a_{t})\mid\pi,P]$. Thus, we can always get $V_{d}({\pi_{k}};P)\leq V_{d}({\pi_{k}};P_{k})+\epsilon_{k}^{\pi_{k}}(P_{k})=V_{\bar{d}}(\pi_{k};P_{k})\leq b_{k}$. Hence, the budget constraint is satisfied for all episodes with high probability. ∎
@@ VI Numerical Results
Figure 3: Cumulative objective gap and cumulative budget violation under (a,b) Poisson arrivals and (c,d) MAWI traces.
@@ VI-A2 Upper level (budget provisioning optimization)
At the beginning of episode $k$, the provisioned budget is $b_{k}\in[B_{0},T]$. The provisioning cost is $f_{k}(b)=M_{1}b+M_{2}(b-\rho_{k})^{2}+M_{0}(b-\rho_{0})^{2}$, where $\rho_{k}=5.0+0.5\sin(2\pi k/2000)+\epsilon_{k}$ $\epsilon_{k}\sim\mathcal{N}(0,0.1^{2})$, $M_{1}=0.01$, $M_{2}=0.05$, and $M_{0}=0.25$. Across episodes we charge the switching cost $\alpha\|b_{k}-b_{k-1}\|^{2}$ with $\alpha=0.5$. We let the scaling parameter $\beta=1.0$.
@@ VI-C Results
We run $K=5\times 10^{4}$ episodes. Fig. 3 compares objective gap and constraint violation. Across both Poisson and MAWI arrivals, our proposed bi-level algorithm achieves the smallest cumulative objective gap while maintaining near-zero violations, demonstrating that dual-coupled provisioning reduces cost without sacrificing safety. In contrast, fixed-budget DOPE incurs substantially larger objective gap due to non-adaptive provisioning, and the fixed-scheduling OCO is consistently worse than our proposed algorithm, indicating that ignoring lower-level sensitivity feedback leads to mis-provisioning under stateful dynamics and switching overhead.

### v2

@@ V-C2 Value approximation error
The remaining terms capture the value difference between the true and surrogate objectives. By definition, $g_{k}(b)-\hat{g}_{k}(b)=\beta(L_{k}^{*}(b)-\hat{L}_{k}(b))$. Using the optimism lemma from the lower-level analysis (Please see Appendix A for complete proof of Lemma 10), we can show that $\hat{L}_{k}(b)$ (the value of the optimistic surrogate) is close to $L_{k}^{*}(b)$ (the true optimal value). The cumulative sum of these differences/errors is bounded by the lower-level regret bound, which thus yields the $\tilde{\mathcal{O}}(\sqrt{K})$ term in the regret. ∎
@@ V-D Final Regret Bound and Safety Guarantee
Theorem 1 establishes a near-optimal $\tilde{\mathcal{O}}(\sqrt{K})$ regret, implying the average regret vanishes. Importantly, the bound separates the roles of the two levels: the upper level contributes only logarithmic regret (and bounded switching overhead) thanks to strong convexity and stable updates, while the dominant $\tilde{\mathcal{O}}(\sqrt{K})$ term comes from learning the unknown lower-level dynamics safely even under time-varying budgets. This clean separation clarifies the benefit of our primal-dual interface. The extended LP provides a dual variable signal that converts lower-level learning progress into correct sensitivity feedback for upper-level subgradient in budget updating. Please see Appendix C for the complete proof of Theorem 1.
@@
Theorem 2 establishes the safety of our BALDE subroutine under dynamic/time-varying budget and constraint threshold. It guarantees that the lower-level scheduler strictly adheres to the resource limits imposed by the upper level. Please see Appendix D for the complete proof of Theorem 2.
@@
For the warm-up phase ($k\leq K_{0}$), safety is guaranteed by baseline policy $\pi_{\mathrm{base}}$. For $k>K_{0}$, applying the value difference lemma and Hölder’s inequality, the difference between the cost under true transition kernel $P$ and that under $P_{k}$ can be bound as $\left|V_{d}({\pi_{k}};P)-V_{d}({\pi_{k}};P_{k})\right|\leq\mathbb{E}[\sum_{t=1}^{T}\|P_{t}-P_{k,t}\|_{1}\|V_{d,t+1}^{\pi_{k}}\|_{\infty}]\leq T\mathbb{E}_{\pi_{k},P_{k}}[\sum_{t=1}^{T}\bar{\beta}_{t,k}^{p}]=\epsilon_{k}^{\pi_{k}}(P_{k})$, where $V_{d,t}^{\pi}\triangleq\mathbb{E}_{\pi,P}[\sum\nolimits_{\tau=t}^{T}d_{t}(s_{t},a_{t})]$. Thus, we can always get $V_{d}({\pi_{k}};P)\leq V_{d}({\pi_{k}};P_{k})+\epsilon_{k}^{\pi_{k}}(P_{k})=V_{\bar{d}}(\pi_{k};P_{k})\leq b_{k}$. Hence, the budget constraint is satisfied for all episodes with high probability. ∎
@@ VI Numerical Results
Figure 3: Cumulative objective gap and cumulative budget violation under (a,b) Poisson arrivals and (c,d) MAWI traces. Here, $\rho_{k}=5.0+0.5\sin(2\pi k/2000)+\epsilon_{k}$, $\epsilon_{k}\sim\mathcal{N}(0,0.1^{2})$, $M_{0}=0.25$, $M_{1}=0.01$, $M_{2}=0.05$, $\alpha=0.5$, $\beta=1.0$.
@@ VI-A2 Upper level (budget provisioning optimization)
At the beginning of episode $k$, the budget $b_{k}\in[B_{0},T]$ is provisioned. The cost $f_{k}(b)=M_{1}b+M_{2}(b-\rho_{k})^{2}+M_{0}(b-\rho_{0})^{2}$ accounts for linear resource expenditures ($M_{1}$) and Service Level Agreement tracking ($M_{2},M_{0}$), where sinusoidal $\rho_{k}$ emulates diurnal seasonality for non-stationary evaluation. Across episodes we charge the switching cost $\alpha\|b_{k}-b_{k-1}\|^{2}$.
@@ VI-C Results
We run $K=5\times 10^{4}$ episodes. Fig. 3 reports the cumulative objective gap (a,c) and cumulative budget violation (b,d) under Poisson arrivals and MAWI traces. In both traffic settings, our bi-level optimization-learning method achieves the smallest objective gap among all budget-feasible algorithms while keeping the cumulative violation essentially zero, indicating that dual-coupled provisioning can reduce long-run cost without sacrificing safety. In contrast, fixed-budget DOPE remains safe but incurs a substantially larger objective gap because the provisioned budget is not adapted across episodes. Finally, the fixed-scheduling OCO baseline attains a smaller objective gap by directly optimizing provisioning costs, but it persistently violates the episode budget, demonstrating that ignoring the cross-level coupling and lower-level safety/sensitivity feedback can lead to systematically infeasible provisioning under stateful queue dynamics and switching overhead.